OpenStudy (anonymous):

A bacteria culture starts with 160 bacteria and grows at a rate proportional to its size. After 4 hours there will be 640 bacteria. (a) Express the population after t hours as a function of t.

OpenStudy (anonymous):

I got y=160 e^(t(In3.75/4))

OpenStudy (anonymous):

) What will be the population after 5 hours? i entered 5 into T

OpenStudy (anonymous):

\[y=Ce^{kt}\] You're told that \(C=160\), and that at \(t=4\) the population is 640. So, you have the equation \[640=160e^{4k}\Rightarrow k=\frac{\ln2}{2}\]

OpenStudy (anonymous):

so my equation was wrong....

OpenStudy (anonymous):

how did you get this..?

OpenStudy (anonymous):

it says i have to write equation : function of t)

OpenStudy (anonymous):

\[640=160e^{4k}\\ 4=e^{4k}\\ \ln4=4k\\ \frac{\ln4}{4}=k\\ \frac{\ln2^2}{4}=k\\ \frac{2\ln2}{4}=k\\ k=\frac{\ln2}{2}\]

OpenStudy (anonymous):

For part a, you need to find the value of k. After that, you'll have both C and k, which you'll replace in the first equation I wrote: \[\large y(t)=160e^{\frac{\ln2}{2}t}\] Then use this to find the population at t = 5, i.e. \(y(5).\)

OpenStudy (anonymous):

wht its become 4=e^4k?

OpenStudy (anonymous):

\(640\div160=4\)

OpenStudy (anonymous):

Oh I see haha the hardest one is the last question c) How long will it take for the population to reach 1290 ? Do i solve for t?

OpenStudy (anonymous):

1290=160e(in2/2)(t)?

OpenStudy (anonymous):

Yeah, that's it. Just make sure you know ln(2)/2 * t is in the exponent.